Calculate the angle of dip if a dip needle oscillating in a vertical plane makes 40 oscillations per min in a magnetic meridian and 30 oscillations per min in a vertical plane at right angle to the magnetic meridian?And is sin@= 0.5625

Grade:12th pass

1 Answers

ROSHAN MUJEEB
3 years ago

We have,
ω
ν

=40/min
ω
H

=30/min
We know that, tanθ=(
ω
ν


ω
H



)
2

tanθ=(
40
30

)
2

θ=tan
−1
(0.75)
2

θ=tan
−1
(0.5625)